OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations.

OpenStudy (anonymous):

can u post the question on here

OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. x - y = 120 x + y = 150

OpenStudy (anonymous):

so use subtraciton x - y = 120 - x + y = 150

jigglypuff314 (jigglypuff314):

let's try to use elimination by addition that is when you look at how it lines up and you add down so like -y + y = 0 which "cancels out" and x + x = 2x and 120 + 150 = 270 so you are left with 2x = 270 do you follow so far?

OpenStudy (anonymous):

yess

jigglypuff314 (jigglypuff314):

then can you solve for x from what I gave? :)

OpenStudy (anonymous):

would it be x=135

jigglypuff314 (jigglypuff314):

correct :) then you can plug that into one of the original equations and solve for y